# Heavy neutral bosons and dark matter in the 3-3-1 model with axionlike particle

### v1

@@ Abstract
We consider heavy neutral bosons in the 3-3-1 model with axionlike particles (331ALP), including the Higgs boson and the $Z^{\prime}$ boson which are outside the standard model (SM). Based on gluon-gluon fusion at the LHC, we investigate the signals of cross-sections in the parameter space region satisfying the current experimental limits of lepton flavor violating decay, including processes involving both charged leptons and Higgs boson, and provide predictions of $m_{h_{2}}\geq 600\penalty 10000\ \mathrm{GeV}$. A new gauge boson, labeled as $Z^{\prime}$, is predicted $m_{Z^{\prime}}\geq 5.1\penalty 10000\ \mathrm{TeV}$ based on the search for high-mass dilepton resonances at ATLAS and CMS. We also introduced the symmetry group $U(1)_{\mathcal{N}}\otimes\mathcal{Z}_{2}^{\prime}$ into the model to indicate dark matter candidates. Investigating the relic density of dark matter within experimentally permissible limits, we established a relationship between the mass of dark matter and the breaking scale of axion.
@@ I Introduction
The LFV processes are of greater interest in the models beyond the standard model (BSM), where there are abundant sources of lepton flavor violation. They may appear alongside mass generation mechanisms for neutrinos, including two main types: effective operators Mizukoshi:2010ky ; Dias:2005yh ; Hung:2022vqx ; Hung:2022kmv and seesaw mechanisms Boucenna:2015zwa ; Hernandez:2013hea ; Nguyen:2018rlb ; Hung:2021fzb ; Ferreira:2019qpf ; CarcamoHernandez:2019pmy ; Catano:2012kw ; Hernandez:2014lpa ; Dias:2012xp . Both these types yielded good results regarding LFVHD, even showing that $Br(h\rightarrow\mu\tau)$ can reach $\mathcal{O}(10^{-4})$ Hue:2015fbb ; Thuc:2016qva ; Zhang:2015csm ; Herrero-Garcia:2017xdu . Furthermore, some publications have indicated a parameter space region that satisfies the experimental limits of both the lepton flavor violating of charged lepton (cLFV) and LFVHD, which is well-suited for investigating other phenomena Herrero-Garcia:2016uab ; Blankenburg:2012ex ; Pilaftsis:1992st .
@@
The paper is organized as follows. In the next section, we review the model and give masses spectrum of gauge and Higgs bosons. We calculate the Feynman rules and analytic formulas for cLFV and LFVHDs in Section III. $Z^{\prime}$ boson and dark mater are discussed in Section.IV. Conclusions are in Section V. Finally, we provide Appendix A,B to calculate the amplitude and show numerical results of $h_{1,2}\rightarrow\mu\tau$ decays.
@@ II.1 Particle content and discrete symmetries
The third component of the $SU(3)_{L}$ triplet leptons are $(\nu_{aR})^{c}$, so the simplest way to give them mass is to use the seesaw mechanism through a combination with $N_{aR}$. To perform this, the $Z_{11}$ group with charged of the form $\omega_{m}\equiv e^{\frac{i2\pi m}{11}},m=0,\pm 1\cdots\pm 5$ is introduced. Furthermore, 3-3-1 models with $\beta=\pm\frac{1}{\sqrt{3}}$ often appear the $SU(3)_{L}$ triplet scalars which have the same quantum numbers, it leads to difficulties in deriving physical states due to self-coupling terms of these fields. To avoid this, the $Z_{2}$ symmetry is introduced into the model, under the transformation of this group, the odd fields are given as, alp331
$(\eta\,,\rho\,,u_{{\color[rgb]{.5,0,.5}a}R}\,,d_{nR},l_{aR},N_{aR})\,\,\,\,\rightarrow\,\,\,\,-\,(\eta\,,\rho\,,u_{aR}\,,d_{nR},l_{aR},N_{aR})\,.$ (4)
@@
$\eta=\left(\begin{array}[]{c}\eta_{1}^{0}\\ \eta_{2}^{-}\\ \eta_{3}^{0}\end{array}\right)\sim\left(1,3,-\frac{1}{3}\right),\,\chi=\left(\begin{array}[]{c}\chi_{1}^{0}\\ \chi_{2}^{-}\\ \chi_{3}^{0}\end{array}\right)\sim\left(1,3,-\frac{1}{3}\right),\,\rho=\left(\begin{array}[]{c}\rho_{1}^{+}\\ \rho_{2}^{0}\\ \rho_{3}^{+}\end{array}\right)\sim\left(1,3,\frac{2}{3}\right),\,\phi\sim(1,1,0)\,.$ (14)
@@
$+\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{3}\left(y_{N}\right)_{ij}\phi\bar{N}_{iR}^{C}N_{jR}+\mbox{H.c.}.$
We can easily see that the Lagrangian in Eq.(25) is conserved charge for all elements of the gauge group, including the $Z_{2}$ group as usual. It is emphasized that the transformation under the $Z_{2}$ as in Eq. (4) is different from the one given in Ref. jpf where $\chi$ is odd. So, $\chi$ only couple with exotic quarks and is responsible for giving mass to particles outside the standard model except $N_{aR}$. The masses of $N_{aR}$ are generated through $v_{\phi}$ of the gauge singlet $\phi$. Furthermore, $v_{\phi}$ also breaks the $Z_{11}$ symmetry to helps the axion-like particle and an inflation arise. Two remaining $SU(3)_{L}$ triplet scalars $\eta$ and $\rho$ generate masses for SM particles as in the usual 3-3-1 models.
@@ II.2 Gauge bosons
where the covariant derivative is as follow,
@@
$P_{\mu}=\frac{g}{2}\left(\begin{array}[]{ccc}W^{3}_{\mu}+\frac{1}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}&\sqrt{2}W^{+}_{\mu}&\sqrt{2}X^{0}_{\mu}\\ \sqrt{2}W^{-}_{\mu}&-W^{3}_{\mu}+\frac{1}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}&\sqrt{2}Y^{-}_{\mu}\\ \sqrt{2}X^{0*}_{\mu}&\sqrt{2}Y^{+}_{\mu}&-\frac{2}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}\\ \end{array}\right),$ (31)
in which we have denoted $t=\frac{g_{X}}{g}$ and defined the mass eigenstates of the charged gauge bosons as
$W^{\pm}_{\mu}$ $=$ $\frac{1}{\sqrt{2}}\left(W^{1}_{\mu}\mp iW^{2}_{\mu}\right),\hskip 8.5359ptY^{\pm}_{\mu}=\frac{1}{\sqrt{2}}\left(W^{6}_{\mu}\pm iW^{7}_{\mu}\right),$
@@
After spontaneous symmetry breaking (SSB), the mass spectrum of the gauge bosons arises from the following kinertic terms:
@@
$m^{2}_{W}=\frac{g^{2}}{4}(v_{\eta}^{2}+v_{\rho}^{2}),\penalty 10000\ m^{2}_{X^{0}}=\frac{g^{2}}{4}(v_{\chi}^{2}+v_{\eta}^{2}),\hskip 8.5359ptm^{2}_{Y}=\frac{g^{2}}{4}(v_{\chi}^{2}+v_{\rho}^{2})\,$ (34)
and their masses splitting with $W$ boson is il :
$|m^{2}_{Y}-m^{2}_{X^{0}}|<m^{2}_{W}\,.$
From (34), it follows
@@
In the basis $(W^{3}_{\mu},W^{8}_{\mu},B_{\mu})$, the squared mass mixing matrix of neutral gauge bosons is:
@@
The matrix $\mathcal{M}^{2}_{NG}$ has three eigenvalues which belong to photon $A$ and $Z$, $Z^{\prime}$ bosons, respectively:
@@ II.3 The scalar sector
The scalar potential of 331ALP has the form:
@@
$+\lambda_{13}(\phi^{*}\phi)(\eta^{\dagger}\eta)+\left(\lambda_{\phi}\epsilon^{ijk}\eta_{i}\rho_{j}\chi_{k}\phi+H.c.\right)$
These scalar fields should be expanded around their VEVs.
@@
The Higgs potential in Eq.(54) gives the mass mixing matrix of the charged scalar fields, diagonalizing it allows to give the physical state and masses of the charged Higgs bosons. The result obtained is similar to Ref.alp331 .
$m^{2}_{H^{\pm}_{1}}=-\frac{v^{2}}{2}\left(\frac{\mathcal{A}}{v_{\rho}^{2}v_{\eta}^{2}}-\lambda_{9}\right)\,,\penalty 10000\ m^{2}_{H^{\pm}_{2}}=-\frac{(v_{\chi}^{2}+v_{\rho}^{2})}{2}\left(\frac{\mathcal{A}}{v_{\chi}^{2}v_{\rho}^{2}}-\lambda_{8}\right)\,.$ (57)
@@
$\left(\begin{array}[]{c}G_{W}^{\pm}\\ H^{\pm}_{1}\\ \end{array}\right)=\left(\begin{array}[]{cc}c_{\alpha}&-s_{\alpha}\\ s_{\alpha}&c_{\alpha}\\ \end{array}\right)\left(\begin{array}[]{c}\rho_{1}^{\pm}\\ \eta_{2}^{\pm}\\ \end{array}\right)\,,\penalty 10000\ \left(\begin{array}[]{c}G_{Y}^{\pm}\\ H^{\pm}_{2}\\ \end{array}\right)=\left(\begin{array}[]{cc}c_{\theta_{1}}&-s_{\theta_{1}}\\ s_{\theta_{1}}&c_{\theta_{1}}\\ \end{array}\right)\left(\begin{array}[]{c}\chi_{2}^{\pm}\\ \rho_{3}^{\pm}\\ \end{array}\right)\,$, (58)
where the mixing angles are defined by $t_{\alpha}=\frac{v_{\eta}}{v_{\rho}}$, $t_{\theta_{1}}=\frac{v_{\rho}}{v_{\chi}}$, and $G_{W}^{\pm}$, $G_{Y}^{\pm}$ are the Goldstone bosons associated with the longitudinal component of the $W^{\pm}$, $Y^{\pm}$, respectively.
@@
Because of the VEVs’ hierachy $v_{\rho},v_{\eta}\ll v_{\chi}\ll v_{\phi}$ as initial assumption, the derived mixing matrix in (67) has just three angles $\alpha,\theta_{3},\theta_{\phi}$ given and one parameter $\left(\frac{1}{v^{2}_{\phi}}+\frac{1}{v^{2}_{\chi}}+\frac{1}{v^{2}_{\rho}}+\frac{1}{v^{2}_{\eta}}\right)$ which is contained in the expression of $A_{5}$ mass in (69). Furthermore, the mass of new massive field CP-odd scalar field $A_{5}$} is given by:
@@
The Eq. ((69)) shows that the values of $\lambda_{\phi}$ should be negative. It is emphasized that both of the squared mass mixing matrix in Eq. (63) and mass of the $A_{5}$ field are arised from the last term in (54) which just appears because of the specific discrete symmetry $Z_{11}\otimes Z_{2}$ imposing to this model from the beginning a2 .
Here, the axionlike particle $a$ is massless and given by the combination of four CP odd neutral scalar fields $I_{\phi}$, $I_{\chi}^{3}$, $I_{\rho}^{2}$ and $I_{\eta}^{1}$ as in (64) which is diffirent from the $a$ field given in Refs.a2 ; jpf . Due to $v_{\chi}\ll v_{\phi}$, the values of $t_{\theta_{\phi}}$ as well as $s_{\theta_{\phi}}$ go to zero, then $c_{\theta_{\phi}}\backsimeq 1$ which approximately helps $a\backsimeq I_{\phi}$.
@@
$H_{\chi}$ $\approx$ $R^{3}_{\chi}c_{\alpha_{\phi}}\,,\mbox{\penalty 10000\ with\penalty 10000\ }m_{H_{\chi}}^{2}{\approx 2\lambda_{1}v_{\chi}^{2}+\frac{\lambda_{5}^{2}}{2\lambda_{1}}v_{\rho}^{2}}\,.$
$\Phi$ $\approx$ $R_{\phi}c_{\alpha_{\phi}}\,,\mbox{\penalty 10000\ with\penalty 10000\ }m_{\Phi}=\sqrt{2\lambda_{10}v_{\phi}}\,$ (82)
In comparision with the $4\times 4$ mass mixing matrix of CP-odd sector containing only four parameters with three massless solution, the matrix in (74) having ten parameters which is unable to be exactly diagonalized. To solve this problem we have used the Hartree-Fock method with conditions such as $v_{\phi}\gg v_{\chi}\gg v_{\rho},v_{\eta}$, $\lambda_{\phi}\ll 1$ and $s_{\alpha_{3}}\approx 0$.
With these VEV hierarchy, the derived matrix contains three angles $\alpha_{2},\alpha_{3}$ and $\alpha_{\phi}$ and three parameters associated with masses of new fields $\Phi,\penalty 10000\ H_{\chi}$ and $h_{2}$. The $\Phi$ field with heavy mass ranging about $10^{11}$ GeV might be used to explain the inflation of the Early Universe alp331 ; Long:2024rdq . The $H_{\chi}$ field is the heavy scalar boson which mass should be at TeV scale. Unfortunately, $\Phi,\penalty 10000\ H_{\chi}$ fields are almost devoid of lepton flavor violating interactions, based on Eqs.25, 75, 82. Only two lighter Higgs bosons, named as $h_{1},\penalty 10000\ h_{2}$, are involved in lepton flavor violating processes, which we will consider below. Both of these particles have masses at electroweak scale, and the lighter particle, $h_{1}$, is identified with the SM-like Higgs boson.

### v2

@@ Abstract
We consider heavy neutral bosons in the 3-3-1 model with axionlike particles (331ALP), including the Higgs boson and the $Z^{\prime}$ boson which are outside the standard model (SM). Based on gluon-gluon fusion at the LHC, we investigate the signals of cross-sections in the parameter space region satisfying the current experimental limits of lepton flavor violating decay, including processes involving both charged leptons and Higgs boson, and provide predictions of $m_{h_{2}}\geq 600\penalty 10000\ \mathrm{GeV}$. A new gauge boson, labeled as $Z^{\prime}$, is predicted $m_{Z^{\prime}}\geq 5.1\penalty 10000\ \mathrm{TeV}$ based on the search for high-mass dilepton resonances at ATLAS and CMS. We consider the stability of odd-$Z_{2}$ particles, with $Z_{2}$ is assumed a residual symmetry after spontaneous symmetry breaking stages, to point out dark matter candidates in the model. Investigating the relic density of dark matter within experimentally permissible limits, we established a relationship between the mass of dark matter and the breaking scale of axion.
@@ I Introduction
The LFV processes are of greater interest in the models beyond the standard model (BSM), where are abundant sources of lepton flavor violation. They may appear alongside mass generation mechanisms for neutrinos, including two main types: effective operators Mizukoshi:2010ky ; Dias:2005yh ; Hung:2022vqx ; Hung:2022kmv ; PhysRevD.47.2918 and seesaw mechanisms Boucenna:2015zwa ; Hernandez:2013hea ; Nguyen:2018rlb ; Hung:2021fzb ; Ferreira:2019qpf ; CarcamoHernandez:2019pmy ; Catano:2012kw ; Hernandez:2014lpa ; Dias:2012xp . Both these types yielded good results regarding LFVHD, even showing that $Br(h\rightarrow\mu\tau)$ can reach $\mathcal{O}(10^{-4})$ Hue:2015fbb ; Thuc:2016qva ; Zhang:2015csm ; Herrero-Garcia:2017xdu . Theoretical results have shown that the signals of LFVHD are much smaller than the experimental limit at the LHC Blankenburg:2012ex , even with the contribution of Majorana neutrinos Pilaftsis:1992st . One publication also pointed out that the LFVHD signal at one loop order is always less than $10^{-4}$, because it is suppressed by a loop factor and constraints from the lepton flavor violating of charged lepton (cLFV) Herrero-Garcia:2016uab . Furthermore, some publications have indicated a parameter space region that satisfies the experimental limits of both cLFV and LFVHD, which is well-suited for investigating other phenomena, for example, explaining the supplemental part for the W boson mass through LFV processes Duy:2024txy , predicting the signal $(g-2)_{\mu}$ in the parameter space region satisfying the experimental limits of LFVHD Hong:2024yhk , studying $Z\rightarrow e_{a}e_{b}$ associated with other LFV processes Hong:2024swk .
@@
The paper is organized as follows. In the next section, we review the model and give masses spectrum of gauge and Higgs bosons. We calculate the Feynman rules and analytic formulas for cLFV and LFVHDs in Section III. $Z^{\prime}$ boson and dark mater are discussed in Section IV and Section V. Conclusions are in Section VI. Finally, we provide Appendix A,B to calculate the amplitude and show numerical results of $h_{1,2}\rightarrow\mu\tau$ decays.
@@ II.1 Particle content and discrete symmetries
The third component of the $SU(3)_{L}$ triplet leptons are $(\nu_{aR})^{c}$, so the simplest way to give them mass is to use the seesaw mechanism through a combination with $N_{aR}$. To perform this, the $Z_{11}$ group with charged of the form $\omega_{m}\equiv e^{\frac{i2\pi m}{11}},m=0,\pm 1\cdots\pm 5$ is introduced. Furthermore, 3-3-1 models with $\beta=\pm\frac{1}{\sqrt{3}}$ often appear the $SU(3)_{L}$ triplet scalars which have the same quantum numbers, it leads to difficulties in deriving physical states due to self-coupling terms of these fields. To avoid this, the $Z_{2}$ symmetry is introduced into the model, under the transformation of this group, the odd fields are given as alp331 ,
$(\eta\,,\rho\,,u_{aR}\,,d_{nR},l_{aR},N_{aR})\,\,\,\,\rightarrow\,\,\,\,-\,(\eta\,,\rho\,,u_{aR}\,,d_{nR},l_{aR},N_{aR})\,.$ (4)
@@
$\eta=\left(\begin{array}[]{c}\eta_{1}^{0}\\ \eta_{2}^{-}\\ \eta_{3}^{0}\end{array}\right)\sim\left(1,3,-\frac{1}{3}\right),\,\chi=\left(\begin{array}[]{c}\chi_{1}^{0}\\ \chi_{2}^{-}\\ \chi_{3}^{0}\end{array}\right)\sim\left(1,3,-\frac{1}{3}\right),\,\rho=\left(\begin{array}[]{c}\rho_{1}^{+}\\ \rho_{2}^{0}\\ \rho_{3}^{+}\end{array}\right)\sim\left(1,3,\frac{2}{3}\right),\,\phi\sim(1,1,0)\,$, (14)
@@
$+\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{3}\left(y_{N}\right)_{ij}\phi\bar{N}_{iR}^{C}N_{jR}+\mbox{H.c.}$
We can easily see that the Lagrangian in Eq.(25) is conserved charge for all elements of the gauge group, including the $Z_{2}$ group as usual. It is emphasized that the transformation under the $Z_{2}$ as in Eq. (4) is different from the one given in Ref. jpf where $\chi$ is odd. So, $\chi$ only couple with exotic quarks and is responsible for giving mass to particles outside the standard model except $N_{aR}$. The masses of $N_{aR}$ are generated through $v_{\phi}$ of the gauge singlet $\phi$. Furthermore, $v_{\phi}$ also breaks the $Z_{11}$ symmetry to helps the axion-like particle and an inflation arise. Two remaining $SU(3)_{L}$ triplet scalars $\eta$ and $\rho$ generate masses for SM particles as in the usual 3-3-1 models. To fully illustrate this, we present the following scheme of symmetry breaking:
$SU(3)_{C}\otimes SU(3)_{L}$ $\otimes$ $U(1)_{X}\otimes Z(11)\otimes Z_{2}$
$\downarrow$ $v_{\phi}$
$SU(3)_{C}\otimes SU(3)_{L}$ $\otimes$ $U(1)_{X}\otimes Z_{2}$
$\downarrow$ $v_{\chi}$
$SU(3)_{C}\otimes SU(2)_{L}$ $\otimes$ $U(1)_{Y}\otimes Z_{2}$
$\downarrow$ $v_{\eta},\,v_{\rho}$
$SU(3)_{C}$ $\otimes$ $U(1)_{Q}\otimes Z_{2}$ (26)
Clearly, $Z_{2}$ is a residual symmetry after the spontaneous symmetry-breaking stages, and it plays a crucial role in identifying stable particles in the model, which may include dark matter candidates.
@@ II.2 Gauge bosons
the covariant derivative is as follow,
@@
$P_{\mu}=\frac{g}{2}\left(\begin{array}[]{ccc}W^{3}_{\mu}+\frac{1}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}&\sqrt{2}W^{+}_{\mu}&\sqrt{2}X^{0}_{\mu}\\ \sqrt{2}W^{-}_{\mu}&-W^{3}_{\mu}+\frac{1}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}&\sqrt{2}V^{-}_{\mu}\\ \sqrt{2}X^{0*}_{\mu}&\sqrt{2}YV^{+}_{\mu}&-\frac{2}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}\\ \end{array}\right),$ (32)
in which we have denoted $t=\frac{g_{X}}{g}$ and defined the mass eigenstates of the charged gauge bosons as:
$W^{\pm}_{\mu}$ $=$ $\frac{1}{\sqrt{2}}\left(W^{1}_{\mu}\mp iW^{2}_{\mu}\right),\hskip 8.5359ptV^{\pm}_{\mu}=\frac{1}{\sqrt{2}}\left(W^{6}_{\mu}\pm iW^{7}_{\mu}\right),$
@@
After spontaneous symmetry breaking (SSB), the mass spectrum of the gauge bosons arises from the following kinertic terms,
@@
$m^{2}_{W}=\frac{g^{2}}{4}(v_{\eta}^{2}+v_{\rho}^{2}),\penalty 10000\ m^{2}_{X^{0}}=\frac{g^{2}}{4}(v_{\chi}^{2}+v_{\eta}^{2}),\hskip 8.5359ptm^{2}_{V}=\frac{g^{2}}{4}(v_{\chi}^{2}+v_{\rho}^{2})\,$ (35)
and their masses splitting with $W$ boson is il
$|m^{2}_{V}-m^{2}_{X^{0}}|<m^{2}_{W}\,.$
From Eq.(35), it follows
@@
In the basis $(W^{3}_{\mu},W^{8}_{\mu},B_{\mu})$, the squared mass mixing matrix of neutral gauge bosons is
@@
The matrix $\mathcal{M}^{2}_{NG}$ has three eigenvalues which belong to photon AA and $Z$, $Z^{\prime}$ bosons, respectively:
@@ II.3 The scalar sector
The scalar potential of 331ALP has the form alp331 :
@@
$+\lambda_{13}(\phi^{*}\phi)(\eta^{\dagger}\eta)+\left(\lambda_{\phi}\epsilon^{ijk}\eta_{i}\rho_{j}\chi_{k}\phi+H.c.\right).$
These scalar fields should be expanded around their VEVs,
@@
The Higgs potential in Eq.(55) gives the mass mixing matrix of the charged scalar fields, diagonalizing it allows to give the physical state and masses of the charged Higgs bosons. The result obtained is similar to Ref.alp331 ,
$m^{2}_{H^{\pm}_{1}}=-\frac{v^{2}}{2}\left(\frac{\mathcal{A}}{v_{\rho}^{2}v_{\eta}^{2}}-\lambda_{9}\right)\,,\penalty 10000\ m^{2}_{H^{\pm}_{2}}=-\frac{(v_{\chi}^{2}+v_{\rho}^{2})}{2}\left(\frac{\mathcal{A}}{v_{\chi}^{2}v_{\rho}^{2}}-\lambda_{8}\right)\,$, (58)
@@
$\left(\begin{array}[]{c}G_{W}^{\pm}\\ H^{\pm}_{1}\\ \end{array}\right)=\left(\begin{array}[]{cc}c_{\alpha}&-s_{\alpha}\\ s_{\alpha}&c_{\alpha}\\ \end{array}\right)\left(\begin{array}[]{c}\rho_{1}^{\pm}\\ \eta_{2}^{\pm}\\ \end{array}\right)\,,\penalty 10000\ \left(\begin{array}[]{c}G_{V}^{\pm}\\ H^{\pm}_{2}\\ \end{array}\right)=\left(\begin{array}[]{cc}c_{\theta_{1}}&-s_{\theta_{1}}\\ s_{\theta_{1}}&c_{\theta_{1}}\\ \end{array}\right)\left(\begin{array}[]{c}\chi_{2}^{\pm}\\ \rho_{3}^{\pm}\\ \end{array}\right)\,$, (59)
where the mixing angles are defined by $t_{\alpha}=\frac{v_{\eta}}{v_{\rho}}$, $t_{\theta_{1}}=\frac{v_{\rho}}{v_{\chi}}$, and $G_{W}^{\pm}$, $G_{V}^{\pm}$ are the Goldstone bosons associated with the longitudinal component of the $W^{\pm}$, $V^{\pm}$, respectively.
@@
Because of the VEVs’ hierachy $v_{\rho},v_{\eta}\ll v_{\chi}\ll v_{\phi}$ as initial assumption, the derived mixing matrix in Eq.(68) has just three angles $\alpha,\theta_{3},\theta_{\phi}$ given and one parameter $\left(\frac{1}{v^{2}_{\phi}}+\frac{1}{v^{2}_{\chi}}+\frac{1}{v^{2}_{\rho}}+\frac{1}{v^{2}_{\eta}}\right)$ which is contained in the expression of $A_{5}$ mass in Eq.(70). Furthermore, the mass of new massive field CP-odd scalar field $A_{5}$} is given by:
@@
The Eq. (70) shows that the values of $\lambda_{\phi}$ should be negative. It is emphasized that both of the squared mass mixing matrix in Eq. (64) and mass of the $A_{5}$ field are arised from the last term in Eq.(55) which just appears because of the specific discrete symmetry $Z_{11}\otimes Z_{2}$ imposing to this model from the beginning a2 .
Here, the axionlike particle $a$ is massless and given by the combination of four CP odd neutral scalar fields $I_{\phi}$, $I_{\chi}^{3}$, $I_{\rho}^{2}$ and $I_{\eta}^{1}$ as in Eq.(65) which is diffirent from the $a$ field given in Refs.a2 ; jpf . Due to $v_{\chi}\ll v_{\phi}$, the values of $t_{\theta_{\phi}}$ as well as $s_{\theta_{\phi}}$ go to zero, then $c_{\theta_{\phi}}\backsimeq 1$ which approximately helps $a\backsimeq I_{\phi}$.
@@
$H_{\chi}$ $\approx$ $R^{3}_{\chi}c_{\alpha_{\phi}}\,,\mbox{\penalty 10000\ with\penalty 10000\ }m_{H_{\chi}}^{2}{\approx 2\lambda_{1}v_{\chi}^{2}+\frac{\lambda_{5}^{2}}{2\lambda_{1}}v_{\rho}^{2}}\,$,
$\Phi$ $\approx$ $R_{\phi}c_{\alpha_{\phi}}\,,\mbox{\penalty 10000\ with\penalty 10000\ }m_{\Phi}=\sqrt{2\lambda_{10}v_{\phi}}.\,$ (83)
In comparision with the $4\times 4$ mass mixing matrix of CP-odd sector containing only four parameters with three massless solution, the matrix in Eq.(75) having ten parameters which is unable to be exactly diagonalized. To solve this problem we have used the Hartree-Fock method with conditions such as $v_{\phi}\gg v_{\chi}\gg v_{\rho},v_{\eta}$, $\lambda_{\phi}\ll 1$ and $s_{\alpha_{3}}\approx 0$.
With these VEV hierarchy, the derived matrix contains three angles $\alpha_{2},\alpha_{3}$ and $\alpha_{\phi}$ and three parameters associated with masses of new fields $\Phi,\penalty 10000\ H_{\chi}$ and $h_{2}$. The $\Phi$ field with heavy mass ranging about $10^{11}$ GeV might be used to explain the inflation of the Early Universe alp331 ; Long:2024rdq . The $H_{\chi}$ field is the heavy scalar boson which mass should be at TeV scale. Unfortunately, $\Phi,\penalty 10000\ H_{\chi}$ fields are almost devoid of lepton flavor violating interactions, based on Eqs.(25, 76, 83). Only two lighter Higgs bosons, named as $h_{1},\penalty 10000\ h_{2}$, are involved in lepton flavor violating processes, which we will consider below. Both of these particles have masses at electroweak scale, and the lighter particle, $h_{1}$, is identified with the SM-like Higgs boson.